OpenStudy (anonymous):

Help. medal to first best answer!:) If the volume of a pyramid is 60 cubic inches and the area of the base is 36 square inches, what is the height of the pyramid?

OpenStudy (mertsj):

\[V=\frac{1}{3}Bh\] where B is the area of the base.

OpenStudy (anonymous):

60=1/3* (36*36)*x 60=1/3*1296*x 60=432x 0.13888...=x

OpenStudy (mertsj):

\[60=\frac{1}{3}(36)h\]

OpenStudy (mertsj):

Please note that 36 is the AREA of the base.

OpenStudy (anonymous):

so if its the area of the base the you dont do 36^2?.... 60=12x 5=x:)

OpenStudy (mertsj):

Remember in the formula B stands for the AREA of the base and 36 is the AREA of the base. So do you think you should square it?

OpenStudy (anonymous):

nope.

OpenStudy (mertsj):

You are correct.

OpenStudy (anonymous):

haha:) Thank you so much!